OpenStudy (anonymous):

Differential Eq: Use Euler's formula to write the given expression in the form a + ib: 1) e^([3pi/4]*i) 2) e^(4-(pi/3)*i)

OpenStudy (anonymous):

\[e^{\frac{3\pi i}{4}}=\cos(\frac{3\pi }{4})+i\sin(\frac{3\pi }{4})\]now it is a matter of evaluating the trig functions at \[\frac{3\pi}{4}\] and that is it. similarly for the other one